Required information The following information applies to the questions displayed below Selected comparative financial statements of Korbin Company follow KORBIN COMPANY Comparative Income Statements For Years Ended December 31, 2017 2016, and 2015 2017 2816 503,920 386,044 267,980 242,822 143,222 53,274 33,972 87,246 2015 Sales Cost of goods sold Gross profit Selling expenses Administrative expenses Total expenses Income before taxes Income taxes Net income 303,360 200,560 171.456 96,444 35,363 71,557 45,353 116,910 22,236 83,658 15,559 11475 55,976 38, 845 7,886 $ 68,091 44,501 38,959 KORBIN COMPANY Comparative Balance Sheets December 31, 2017, 2016, and 2015 2016 51,214 40,069 53,563 95,112 101,019 60,047 2017 2015 Assets Current assets Long-term investments Plant assets, net Total assets 1,089 4,110 $140,326 142,088 117,720 Liabilities and Equity Current liabilities Common stock Other paid-in capital Retained earnings Total liabilities and equity s 21,364 $ 21,171 20,601 65,000 65,000 8,125 47,000 5,222 47,79244 897 8,125 $146,326 142,888 117,720 2. Complete the below table to calculate income statement data in common-size percents. (Round your percentage answers to 2 decimal places.) KORBIN COMPANY Common Size Comparative Income Statements
2 Complete the below table to calculate income statement data in common-size percents (Round your percentage answers to 2 decimal places. KORBIN COMPANY Common-Size Comparative Income Statements For Years Ended December 31, 2017, 2016, and 2015 2017201 2015 Sales Cost of goods sold Gross proft Selling expenses Administrative expenses Total expenses Income before taxes Income taxes Net income
Answer #1

wNyyxAFFX0UdgAAAABJRU5ErkJggg==

% is calculated dividing the amount by sales amount

Example

sales/Sales =503920/503920=1

Cost of Goods sold =303360/503920=60.20%